jess chris and nick scored a total of 249 points on their math exam. jess score was 20 points higher than chris's score. twice nicks score was 6 more than the sum of chris's and jess's score waht was each of their scores on the exam?
Answer by stanbon(75887)

jess chris and nick scored a total of 249 points on their math exam. jess score was 20 points higher than chris's score. twice nicks score was 6 more than the sum of chris's and jess's score waht was each of their scores on the exam?
-------------------------
Equations:
j + c + n = 249 points
----
j = c+20
2n = c + j + 6
---------------------------
Modify by substituting:
2n = c + c+20 + 6
2n = 2c + 26
n = c + 13
----
Substitute for j and n and solve for "c":
c+20 + c+13 + c = 249
3c + 33 = 249
3c = 216
c = 72 (# of points Chris has)
j = c + 20 = 92 (# of points jess has)
n = c + 13 = 85 (# of points nick has)
========================
Cheers,
Stan H.